Area of five square of the same size is 245cm then its perimeter will
Answer by dkppathak(439)

Area of five square of the same size is 245cm then its perimeter will
area of one square is x^2
than are of 5 such square is 5x^2
5x^2=245
x^2=245/5=49
X=7 cm
length of five such square will be 5x7 =35 cm
breath will be 7 cm only
perimeter =2(L+B)= 2(35+7)=2x42=84 cm